# Localisation of Regularised and Multiview Support Vector Machine Learning

### v3

@@ 3. Differentiable Loss Functions
Theorem 3.7 and its Corollary 3.9 have a theoretical significance and a less practical importance because, in the system (3.4) and (3.5), the unknowns $a_{1},\ldots,a_{l+u}$ appear also on the right hand side and hence it is more an implicit form of expressing them and not an explicit one. Because of that, for specified loss functions $V$, one should work further on these expressions in order to obtain explicit, or at least computable, solutions. In the next two subsections, we work out the details and show how Theorem 3.7 and its Corollary 3.9 can be improved for special loss functions, the least squares functions, similar to the results in [48], and the exponential least squares functions, and compare the formulae.
@@
$a_{i}$ $+\sum_{j=1}^{l+u}\Bigl{[}\frac{\exp(-\|y_{j}-C_{x_{j}}\sum\limits_{k=1}^{l+u}K(x_{j},x_{k})a_{k}\|^{2}_{{\mathcal{Y}}_{x_{j}}})}{l\gamma_{A}}C_{x_{i}}^{*}C_{x_{i}}K(x_{i},x_{j})+\frac{\gamma_{I}}{\gamma_{A}}\sum_{k=1}^{l+u}M_{i,k}K(x_{k},x_{j})\Bigl{]}a_{j}$
@@
3.5. An Algorithm for the Exponential Least Square Loss Function.
@@
where
@@
$H_{i}(\boldsymbol{a})$ $=a_{i}+\frac{1}{l\gamma_{A}}\sum_{j=1}^{l+u}\exp\bigl{(}-\|y_{j}-C_{x_{j}}\sum\limits_{k=1}^{l+u}K(x_{j},x_{k})a_{k}\|^{2}_{{\mathcal{Y}}_{x_{j}}}\bigr{)}C_{x_{i}}^{*}C_{x_{i}}K(x_{i},x_{j})a_{j}$
@@
Here we show how the potential reduction Newton’s method in [50] can be used in order to provide an algorithm to approximate solutions $\boldsymbol{a}\in{\mathbb{R}}^{N}$ of the equation $H(\boldsymbol{a})=0$ under the assumption that the Jacobi matrix $\nabla_{\boldsymbol{u}}H$ is nonsingular for all $\boldsymbol{u}\in{\mathbb{R}}^{N}$. Following [50] we consider the potential function $p\colon{\mathbb{R}}^{N}\rightarrow{\mathbb{R}}$ given by $p(\boldsymbol{u})=\|\boldsymbol{u}\|^{2}$, $\boldsymbol{u}\in{\mathbb{R}}^{N}$. Note that $p$ is continuously differentiable, that $\boldsymbol{u}^{T}\nabla_{\boldsymbol{u}}p>0$ for all nonzero $\boldsymbol{u}\in{\mathbb{R}}^{N}$ and that $\lim_{\|\boldsymbol{u}\|\rightarrow+\infty}p(\boldsymbol{u})=+\infty$. In addition, for any $\bar{\sigma}\in(0,1]$ and any $\boldsymbol{b}\in{\mathbb{R}}^{N}$ we have
$\bar{\sigma}(\boldsymbol{b}^{T}\boldsymbol{u})(\boldsymbol{b}^{T}\nabla_{\boldsymbol{u}}p)\leq\|\boldsymbol{b}\|^{2}(\boldsymbol{u}^{T}\nabla_{\boldsymbol{u}}p),\quad\boldsymbol{u}\in{\mathbb{R}}^{N}.$
Indeed, since $p(\boldsymbol{u})=\|\boldsymbol{u}\|^{2}$ this is a consequence of the Cauchy inequality.
The previous considerations show that we can obtain the following algorithm for approximation of solutions of the equation $H(\boldsymbol{a})=0$ derived from the unconstrained version of the potential reduction Newton’s method described in [50]. Fix $\bar{\sigma}\in(0,1]$ and $\boldsymbol{b}\in{\mathbb{R}}^{N}$.
Step 0. (Initialisation) Let a vector $\boldsymbol{a}_{0}\in{\mathbb{R}}^{N}$, scalars $\rho\in(0,1)$ and $\alpha\in(0,1)$, a numerical sequence $(\sigma_{n})_{n}$, with $\sigma_{n}\in(0,\bar{\sigma}]$, and the prescribed tolerance $\epsilon>0$ be given. The iteration counter is set to $n=0.$
Step 1. (Computing the modified Newton direction) Calculate the vector $\boldsymbol{d}_{n}\in{\mathbb{R}}^{N}$ which is the solutione $\boldsymbol{d}$ of the linear equation
$H(\boldsymbol{a}_{n})+(\nabla_{a_{n}}H)\boldsymbol{d}=\frac{\sigma_{n}\boldsymbol{b}^{T}H(\boldsymbol{a}_{n})}{\|\boldsymbol{b}\|^{2}}\boldsymbol{b}.$ (3.31)
This equation has unique solution $\boldsymbol{d}$ due to the assumption on the nonsingularity of $\nabla_{\boldsymbol{u}}H$ for all $\boldsymbol{u}\in{\mathbb{R}}^{N}$. N.B. If $\boldsymbol{d}$ is the zero vector, the right-hand side of the previous equation is by convention set to 00.
Step 2. (Armijo line search) Let $m_{n}$ be the smallest nonnegative integer $m$ which satisfies the inequality
$\|H(\boldsymbol{a}_{n}+\rho^{m}\boldsymbol{d}_{n})\|^{2}-\|H(\boldsymbol{a}_{n})\|^{2}\leq\alpha\rho^{m}(\nabla_{\boldsymbol{a}_{n}}\|H\|^{2})^{T}\boldsymbol{d}_{n}.$ (3.32)
The existence of $m_{n}$ is guaranteed by the Lagrange Mean Value Theorem and the choice of $\boldsymbol{d}_{n}$ as in (3.31). Then set
$\boldsymbol{a}_{n+1}=\boldsymbol{a}_{n}+\rho^{m_{n}}\boldsymbol{d}_{n}.$ (3.33)
Step 3. (Termination test) If
$\|H(\boldsymbol{a}_{n+1})\|\leq\epsilon$
we stop and accept $\boldsymbol{a}_{n+1}$ as an approximate solution of the equation $H(\boldsymbol{a})=0$. Otherwise, return to Step 1 and replace $n$ by $n+1$.
Remark 3.10.
@@
Because the optimisation problem is nonlinear, multiple solutions may show up and this makes the choice of the initialisation vector $\boldsymbol{a}_{0}$ very important: for different choices of the initialisation vector different pools of solutions of the equations $H(\boldsymbol{a})=0$ might be found. Let us observe that, on the one hand, the components $H_{i}$ for $i=l+1,\ldots,l+u$, corresponding to unlabeled points $x_{i}$, are linear and homogeneous and hence that 00 is a solution. On the other hand, the components $H_{i}$ for $i=1,\ldots,l$, corresponding to labeled points $x_{i}$, are nonlinear and nonhomogeneous and hence that 00 is not a solution. From here, we can see that, on the one hand, as the learning problem is semisupervised, $l$ is significantly less than $u$ and hence taking the initialisation vector $\boldsymbol{a}_{0}=0$ might be a good choice for the beginning. On the other hand, in order to find better minimisers, or even the global minimiser, some multigrid methods or stochastic methods for the choice of the initialisation vector $\boldsymbol{a}_{0}$ might be involved, e.g. see [26].
As a particular case of Theorem 3 in [50] we have the following convergence result.
Proposition 3.11.
With notation as before, assume that the Jacobi matrix $\nabla_{\boldsymbol{u}}H$ is nonsingular for all $\boldsymbol{u}\in{\mathbb{R}}^{N}$ and that $\limsup_{n\rightarrow\infty}\sigma_{n}<\bar{\sigma}$. Then, the sequence $(\boldsymbol{a}_{n})_{n}$ generated by the unconstrained potential reduction Newton algorithm described before is bounded and $\lim_{n\rightarrow\infty}H(\boldsymbol{a}_{n})=0$. In particular, any accumulation point $\boldsymbol{a}$ of the set $\{\boldsymbol{a}_{n}\mid n\in{\mathbb{N}}\}$ is a solution of the equation $H(\boldsymbol{a})=0$.
@@
Because there is no guarantee for uniqueness of solution, this algorithm should be complemented by further steps in which the obtained solutions should be tested whether they provide minimisers or not and to which extent from the class of all minimisers one can get a global minimiser. But the most challenging problem refers to the assumption that the Jacobi matrix $\nabla_{\boldsymbol{u}}H$ is nonsingular for all $\boldsymbol{u}\in{\mathbb{R}}^{N}$. In order to tackle this question we explicitly calculate this Jacobi matrix. To this end, we first observe that for each $i,j=1,\ldots,l+u$, letting $a_{j}=(a_{j}^{(1)},\dots,a_{j}^{(d_{j})})$, we have a partial Jacobi matrix
@@
$\frac{\partial H_{i}}{\partial a_{i}}$ $=I_{{\mathcal{W}}_{x_{i}}}+\frac{\gamma_{I}}{\gamma_{A}}\sum_{k=1}^{l+u}M_{i,k}K(x_{k},x_{i}),$ (3.36)

### v4

@@ 3. Differentiable Loss Functions
Theorem 3.7 and its Corollary 3.9 have a theoretical significance and a less practical importance because, in the system (3.4) and (3.5), the unknowns $a_{1},\ldots,a_{l+u}$ appear also on the right hand side and hence it is more an implicit form of expressing them and not an explicit one. Because of that, for specified loss functions $V$, one should work further on these expressions in order to obtain explicit or, at least computable, solutions. In the next two subsections, we work out the details and show how Theorem 3.7 and its Corollary 3.9 can be improved for special loss functions, the least squares functions, similar to the results in [51], and the exponential least squares functions, and compare the formulae.
@@
$a_{i}$ $+\sum_{j=1}^{l}\Bigl{[}\frac{\exp(-\|y_{j}-C_{x_{j}}\sum\limits_{k=1}^{l+u}K(x_{j},x_{k})a_{k}\|^{2}_{{\mathcal{Y}}_{x_{j}}})}{l\gamma_{A}}C_{x_{i}}^{*}C_{x_{i}}K(x_{i},x_{j})+\frac{\gamma_{I}}{\gamma_{A}}\sum_{k=1}^{l+u}M_{i,k}K(x_{k},x_{j})\Bigl{]}a_{j}$
@@
3.5. Available Numerical Methods for the Exponential Least Square Loss Function.
@@
where $a_{j}\in{\mathcal{W}}_{x_{j}}={\mathbb{R}}^{d_{j}}$, for each $j=1,\ldots,l+u$, and
@@
$H_{i}(\boldsymbol{a})$ $=a_{i}+\frac{1}{l\gamma_{A}}\sum_{j=1}^{l}\exp\bigl{(}-\|y_{j}-C_{x_{j}}\sum\limits_{k=1}^{l+u}K(x_{j},x_{k})a_{k}\|^{2}_{{\mathcal{Y}}_{x_{j}}}\bigr{)}C_{x_{i}}^{*}C_{x_{i}}K(x_{i},x_{j})a_{j}$
@@
There are different methods and algorithms for dealing with constrained nonlinear systems of differentiable functions. For example, the potential reduction Newton’s method in [53] can be used in order to provide an algorithm to approximate solutions $\boldsymbol{a}\in{\mathbb{R}}^{N}$ of the equation $H(\boldsymbol{a})=0$ under the assumption that the Jacobi matrix $\nabla_{\boldsymbol{u}}H$ is nonsingular for all $\boldsymbol{u}\in{\mathbb{R}}^{N}$. Other methods based on the interior point methods are available, for example see [15], [16], and [70]. The latter algorithm is implemented in the fmincon function in MATLAB, that we have used in our example.
@@
Because the optimisation problem is nonlinear, multiple solutions may show up and this makes the choice of the initialisation vector $\boldsymbol{a}_{0}$ very important: for different choices of the initialisation vector different pools of solutions of the equations $H(\boldsymbol{a})=0$ might be found. Let us observe that, on the one hand, the components $H_{i}$ for $i=l+1,\ldots,l+u$, corresponding to unlabeled points $x_{i}$, are linear and homogeneous and hence that 00 is a solution. On the other hand, the components $H_{i}$ for $i=1,\ldots,l$, corresponding to labeled points $x_{i}$, are nonlinear and nonhomogeneous and hence that 00 is not a solution. From here, we can see that, on the one hand, as the learning problem is semisupervised, $l$ is significantly less than $u$ and hence taking the initialisation vector $\boldsymbol{a}_{0}=0$ might be a good choice for the beginning. On the other hand, in order to find better minimisers, or even the global minimiser, some multigrid methods or stochastic methods for the choice of the initialisation vector $\boldsymbol{a}_{0}$ might be involved, e.g. see [28]. In our example, we have used the Latin Hypercube Sampling (LHS), see [46], to solve this issue. Similarly to multigrid methods, the LHS divides the domain into small cubes. Without loss of generality, assume that the domain is $[0,1]^{d}$ a cube with dimension $d$. Then for a fixed $n$, split the cube into $n^{d}$ subcubes by splitting each interval $[0,1]$ into intervals with $1/n$ length, that is,
$[0,1]=\bigcup_{i=0}^{n-1}\left[\dfrac{i}{n},\dfrac{i+1}{n}\right].$
Then we pick points such that any rectangle $R_{i,j}$ defined below contains only one point
$R_{i,j}=[0,1]^{j-1}\times\left[\dfrac{i}{n},\dfrac{i+1}{n}\right]\times[0,1]^{d-j},\quad i,j=1,\dots,n.$
This sampling strategy enjoys asymptotic bounds in expectation [65] with few points. Hence it is a good candidate for our purposes.
@@
Because there is no guarantee for uniqueness of solution, the algorithms for approximation of the solutions of the equation $H(\boldsymbol{a})=0$ should be complemented by further steps in which the obtained solutions should be tested whether they provide minimisers, of the learning function (3.45), or not and to which extent from the class of all minimisers one can get a global minimiser. But the most challenging problem refers to the assumption that the Jacobi matrix $\nabla_{\boldsymbol{u}}H$ is nonsingular for all $\boldsymbol{u}\in{\mathbb{R}}^{N}$. In order to tackle this question we explicitly calculate this Jacobi matrix. To this end, we first observe that for each $i,j=1,\ldots,l+u$, letting $a_{j}=(a_{j}^{(1)},\dots,a_{j}^{(d_{j})})$, we have a partial Jacobi matrix
@@
$\frac{\partial H_{i}}{\partial a_{i}}$ $=I_{d_{i}}+\frac{\gamma_{I}}{\gamma_{A}}\sum_{k=1}^{l+u}M_{i,k}K(x_{k},x_{i}),$ (3.33)